OpenStudy (taepoohwae):

use the elimination method to solve. a+7b=13 a+2b=-2 Is there a solution Are there infinetly many soutions Or is there no solution

OpenStudy (anonymous):

did you try the method of elimination?

OpenStudy (anonymous):

ubtract the bottom equation from the top equation, then solve for b and then plug the value you get back into either equation

OpenStudy (anonymous):

if you try the method of elimination and it works for finding the unknown, then there are two ways in doing so

OpenStudy (anonymous):

Let's eliminate a since the coefficient of a is the same in both equation. a + 7b = 13 -a + 2b = -2 ------------- 5b = 15 b = 3 Plug in b = 3 in to one of the equations and solve for a: a + 7(3) = 13 a = 13 - 21 a = -8 Therefore, b = 3 and a = -8

OpenStudy (taepoohwae):

Thanks genius12! just beginning algebra1 at the age of 40 and have no clue as to what i'm doing. I'll use this as notes, it looks fairly easy.

OpenStudy (anonymous):

Medal? @taepoohwae

OpenStudy (anonymous):

tae, just so you don't get confused what he wrote should be written as a+7b=13 -(a+2b)= -(-2) which becomes -a-2b = 2

OpenStudy (taepoohwae):

Thanks looking through my textbook I've noticed those also.